# Subtree Mode and Applications

### v1

@@ 5 String-Processing Applications
Using Theorem 2 on $\textsf{ST}(S)$, we can solve the Bottom-11 DR problem, which analogously to Bottom $k$-DR [NT15], asks for the string from $\mathcal{S}$ that contains the least number of occurrences of $P$, within the complexities of Theorem 3.
@@
Given a query with $P\in\Sigma^{m}$, we construct the suffix tree $\textsf{ST}(P)$ using Lemma 1, and compute $|\textsf{occ}_{P}(Q)|$, for every string $Q$ of length $q$ that occurs in $P$ in $\mathcal{O}(m)$ total time. We also use $\textsf{ST}(P)$ to mark every position $i^{\prime}$ of $P$ such that $P[i\mathinner{.\,.}i+q)=P[i^{\prime}\mathinner{.\,.}i^{\prime}+q)$, for some $i<i^{\prime}$, to avoid double-counting. We then spell $P$ in $\textsf{ST}(S)$ using suffix links [Gus97], to iterate, for increasing $i$, over the loci (in $\textsf{ST}(S)$) of the longest prefixes of $P[i\mathinner{.\,.}i+q)$ that occur in $S$. We maintain a counter, which is initialized as zero, as follows. If $Q:=P[i\mathinner{.\,.}i+q)$ occurs in $\mathcal{S}$, $i$ is not marked, and $|\textsf{occ}_{P}(Q)|\in[f^{\min}_{v}-\varepsilon,f^{\max}_{v}+\varepsilon]$}, where $v$ is the node of $\textsf{ST}(S)$ with path-label $Q$ (if this is an implicit node, we take its nearest explicit descendant as $v$), we increase the counter by one. The value of the counter is output after processing all of $P$. Spelling $P$ takes $\mathcal{O}(m)$ time [Gus97]; for every $q$-gram $Q$, we make $\mathcal{O}(1)$ elementary $\mathcal{O}(1)$-time operations, and hence the query time follows. ∎
@@ 6 Generalizations of SM
The $k$-RM problem asks for the $k$ most frequent colors in a range of an array [Cha+14]. Analogously, we define the $k$-SM problem asking for the $k$ most frequent colors in the leaves of the subtree rooted at a node of $\mathcal{T}$ (breaking ties arbitrarily). Corollary 2 is obtained by slightly modifying the SCM algorithm from Section 4.
@@
The SCM algorithm naturally extends for $k$-SM. For every node $v$ of $\mathcal{T}$, in Step 3, we now store the $k$ most frequent colors (breaking ties arbitrarily) and their frequencies as a list $(c^{1}_{v},f^{1}_{v}),\dots,(c^{k}_{v},f^{k}_{v})$. (If, for some node, we have fewer than $k$ colors, we leave some entries undefined.) We thus only need to verify that given $n>k$ integers we can select the $k$ largest ones in $\mathcal{O}(n)$ time to obtain the claimed generalization. We do this using the classic linear-time selection algorithm [Blu+73]. After computing one list $(c_{1},f_{1}),\dots,(c_{k},f_{k})$ per node, we sort the lists by frequency using a single global radix sort. This takes $\mathcal{O}(kN)$ time since each frequency is at most $N$. ∎
@@ 7 Lower Bounds for Descendant-Mode in DAGs
Theorem 7 implies that an algorithm for DM that is similarly fast to our SCM algorithm is highly unlikely, as $p(N)=\mathcal{O}(N)$ and $q(N)=\mathcal{O}(1)$ for SM, and this would contradict the combinatorial BMM conjecture [Abb+24]. More generally, for any $\varepsilon>0$, there is no data structure for DM with construction time $\mathcal{O}(N^{\omega/2-\varepsilon})$ and sub-polynomial query time (unless the square matrix multiplication exponent $\omega$ is $2$). Under the combinatorial BMM conjecture, for any $\varepsilon_{1},\varepsilon_{2}\geq 0$, there is no combinatorial data structure for DM with construction time $\mathcal{O}(N^{3/2-\varepsilon_{1}})$ and query time $\mathcal{O}(N^{1/2-\varepsilon_{2}})$. These conditional lower bounds are identical to the ones for RM proved by Chan et al. [Cha+14].
@@ 8 Related Work
Krizanc et al. [KMS05] proposed a data structure for RM with $\mathcal{O}(1)$-time queries and $\mathcal{O}(N^{2}\log\log N/\log N)$ space, and another with $\mathcal{O}(\sqrt{N}\log N)$-time queries and $\mathcal{O}(N)$ space. Chan et al. [Cha+14] improved the time/space trade-off of the data structures of [KMS05] by designing an $\mathcal{O}(N+s^{2}/w)$-space data structure with $\mathcal{O}(N/s)$ query time, for any $s\in[1,N]$, where $w=\Omega(\log N)$ is the machine word size [Cha+14, Section 6]. By setting $s:=\lceil\sqrt{Nw}\rceil$, said data structure takes $\mathcal{O}(N)$ space and has $\mathcal{O}(\sqrt{N/w})$-time queries. In Baseline 2, we employ the data structure in [Cha+14, Section 3], as it is less expensive to construct [Cha+14, Kar+24] and thus leads to a faster baseline for SM (Proposition 2). Greve et al. [Gre+10] showed a lower bound for RM: any data structure that uses $N\log^{\mathcal{O}(1)}N$ space needs $\Omega(\log N/\log\log N)$ time to answer an RM query, and any data structure that supports RM queries in $\mathcal{O}(1)$ time needs $N^{1+\Omega(1)}$ space. The current best conditional lower bound [Cha+14] indicates that answering $N$ RM queries on an array of size $\mathcal{O}(N)$ cannot be performed in $\mathcal{O}(N^{\omega/2-\varepsilon})$ time for any $\varepsilon>0$, where $\omega<2.3716$ [Vas+24] is the square matrix multiplication exponent. Dynamic and multidimensional versions of RM were studied in [SX20, El-+18] and [Dur+15], respectively. RM on paths of a tree (instead of subtrees as in SM) was studied in [Dur+16, GH22].
@@
SM can be used to construct a data structure for the $1$-DR problem; see Section 5. The $k$-DR problem has been studied in theoretical computer science; [NN12, Hon+14, NN17, NN25] proposed linear-space data structures with optimal or near-optimal query time but no efficient algorithms to construct them. $k$-DR differs from Top-$k$ document retrieval based on known patterns (keywords), for which we refer to [KHE20, Gou+25].
@@
The UPM problem in Section 5 falls into the area of pattern mining [Agg14]. Specifically, it is somewhat related to discriminative (a.k.a emerging or contrast sets) pattern mining [DL99, DL05, BP01]. Given a collection of records (sequences [Liu+15, Mat+21, Cha+03], transactions [DL05], relational tuples [DL05], or vectors [BP01]), the latter problem asks for mining all patterns (subsequences [Liu+15, Mat+21], substrings [Cha+03], itemsets [DL05], sets of relational attribute values [DL05], or sets of attribute/value pairs [BP01]) which occur “disproportionately” in two or more parts of the collection that have different class labels. The disproportionality is measured based on difference in frequency [BP01], growth rate [DL99, DL05], weighted relative accuracy [Mat+21], or other measures [Liu+15]. The UPM problem is also relevant to fair pattern mining [Haj+14]. The latter problem asks for mining itemsets in a transaction dataset that do not produce association rules which are unprotected according to legally-grounded fairness measures [Haj+14]. None of the aforementioned approaches can solve UPM.
@@
Sequence comparison by means of $q$-grams is ubiquitous in bioinformatics [Ukk92]. It offers a faster alternative to using more expensive string measures such as edit distance. For instance, BLAST [Alt+90] and FASTA [PL88], two of the most widely-used tools for sequence-to-database search, are based on the notion of $q$-grams to report the best hits for a query.
@@
There are many other problems on leaf-colored [Hui92, Ste93, Mut02] and node-colored [Gaw+18, MS07, FG08] trees.
@@ 9 Experimental Evaluation
We used $4$ benchmark datasets (see Table 2): (1) WebKB [Cra+98], which is a collection of webpages of computer science departments of various universities; (2) Genes, which is a collection of DNA sequences between two markers flanking the human X chromosome centromere [Say+23]; (3) News [ZCG15], which is a collection of documents from the Newsgroups dataset, and (4) Vir [Nat25a], which is a collection of viral genomes. As the baselines could not run on large datasets, we applied them to samples of the first two datasets, constructed by selecting strings that have the same length uniformly at random; see Table 2.
@@
WebKB [Cra+98] Web 26 790,340 32.98 26,068,175 38,491,476
Genes [Nat13] Biology 4 800,000 1,250.01 1,000,007,888 2,030,921,400
News [ZCG15] News 26 4,781 725.556 3,768,883 5,293,347
Vir [Nat25a] Biology 4 143,588 29,085 4,176,208,246 7,313,326,212
@@
WebKB-sam [Cra+98] Web 26 31,030 100 3,103,000 4,876,306
Genes-sam [Nat13] Biology 4 10,000 200 2,000,000 3,753,134
@@
We examined the impact of the two problem parameters, $N$ and $\Delta$, on runtime and space. We also showcase the benefit of our approach in the UPM and the CQS applications. In these applications, we used four real datasets. We did not examine the $1$-DR application, as the existing approaches for $k$-DR do not show how their data structures can be constructed efficiently; and then the querying part of our method reduces to the standard pattern matching on the suffix tree.
@@
Figs. 6(a) to 6(c) show the runtime for varying $N$ and fixed $\Delta$. Our SCM algorithm was faster than the fastest baseline, BA3, by at least one order of magnitude and $24$ times on average. BA3 in turn was faster than both BA1 and BA2 by at least two orders of magnitude on average. All algorithms scaled linearly with $N$, in line with their time complexities, except BA2 whose time complexity is $\mathcal{O}(N\sqrt{N})$. In fact, since the space complexity of BA2 is also $\mathcal{O}(N\sqrt{N})$, it did not terminate for strings with more than 1 million letters; it needed more than the 1TB of memory that was available. Figs. 6(d) to 6(f) show the peak memory consumption for the experiments of Figs. 6(a) to 6(c). SCM needed on average $26\%$ and up to $58\%$ less memory compared to the best baseline, BA3, as storing counts for the single-color trees needs less memory than merging trees. BA3 in turn needed two orders of magnitude less memory on average compared to both BA1 and BA2, as its space complexity is $\mathcal{O}(N)$, while that of BA1 and BA2 is $\mathcal{O}(N\Delta)$ and $\mathcal{O}(N\sqrt{N})$, respectively.
@@
Figs. 7(a) to 7(c) show the runtime for varying $\Delta$ and fixed $N$; we fixed $N=300,000$ for the WebKB-sam dataset, $N=162,000$ for News, and $N=200,000$ for Genes-sam by taking the prefix of length $\lfloor N/\Delta\rfloor$ of each string. Our SCM algorithm was faster than the fastest baseline, BA3, by at least an order of magnitude and $21$ times on average and, as expected by its time complexity, its runtime was not affected by $\Delta$. BA3 in turn was faster than both BA1 and BA2 by two orders of magnitude on average, which is in line with the time complexities of these algorithms. Figs. 7(d) to 7(f) show the peak memory consumption for the experiments of Figs. 7(a) to 7(c). SCM needed at least $20\%$ and up to $50\%$ less memory compared to the best baseline, BA3, for the same reason as in the experiments of Figs. 6(d) to 6(f). BA3 in turn was more space-efficient than both BA1 and BA2 by more than two orders of magnitude on average, which is in line with the space complexities of these algorithms. Also, BA2 uses more memory as $\Delta$ increases, as the string index it uses gets larger.
@@
Figs. 8(a) to 8(c) show the runtime for varying $N$ and fixed $\Delta$. Both SCM and BA3 scaled linearly with $N$, in line with their time complexities, but SCM was at least $18$ and up to $40$ times faster, as BA3 has an extra $\log\Delta$ term in its time complexity which is $17$ to $20$ depending on the dataset. SCM is practical; it took less than $20$ minutes when applied to the entire Vir dataset whose total length is $4.2$ billion letters (and suffix tree has over $7.3$ billion nodes). Figs. 8(d) to 8(f) show the peak memory consumption for the experiments of Figs. 8(a) to 8(c) . SCM needed $15\%$ less memory compared to BA3 on average, which is in line with the space complexities of the algorithms. This shows again the benefit of the count merging in SCM compared to the tree merging in BA3.
@@
Figs. 8(g) to 8(i) show the runtime for varying $\Delta$ and fixed $N$; we fixed $N$ by removing letters from each string evenly, so that the remaining strings have total length $N=2,000,000$ for WebKB, $N=62,500,000$ for Genes, and $N=10^{9}$ for Vir. Again, SCM was faster than BA3 by at least $15$ and $26$ times on average. For example, when $\Delta=800,000$ in Genes (see Fig. 8(h)), SCM took only about $20$ seconds while BA3 about $8$ minutes. Figs. 8(j) to 8(l) show the peak memory consumption for the experiments of Figs. 8(g) to 8(i). Again, SCM was more space-efficient than BA3; it needed at least $9\%$ and up to $17\%$ less memory.
@@
We consider collections whose strings represent different user subpopulations and solve the UPM problem with large $\varepsilon$. We identify patterns revealing behavioral preferences that prevail in these subpopulations. For each pattern, we refer to literature supporting that it is in fact prevailing in these subpopulations.
@@
We processed three datasets: Movielens [HK15], Book-Crossing [Zie+05], and Alibaba [Pei+19]. The processed datasets are comprised of two strings each, and they can be found in https://edu.nl/44ua9. Each string corresponds to a different user subpopulation. In Movielens, one string corresponds to $1,709$ men and the other to $1,709$ women. Each string has length $4,000,033$ and contains genres of rated movies, ordered chronologically. The total number of distinct genres (alphabet size) is $18$. In Book-Crossing, one string corresponds to $9,164$ teenagers (1010-$18$ years old) and the other to $4,252$ elderly (6565-$100$ years old). Each string has length $14,145$ and contains book genres, ordered by users’ ratings for book genres from high to low. The genres were found by mapping book ISBNs to genres using ChatGPT-3.5-turbo [Ope23], which is remarkably good in this task [Raj+25]. The total number of distinct book genres (alphabet size) is $10$. In Alibaba, one string corresponds to $928,622$ users with high purchase power ($\geq 15$) and the other to $894,770$ users with low purchase power (<15<15). Each string has length $34,651,424$ and contains price tiers of products, ordered in the way users browsed them. The total number of distinct price tiers (alphabet size) is $4$. Specifically, the letters L, M, H, and P, correspond to the Low, Middle, High, and Premium price tier, respectively.
@@
Movielens Action Drama War Action 1,551 (Men) 588 (Women) 1,000 [WLS17, KK05]
Movielens Comedy Drama Romance Comedy 899 (Men) 1,860 (Women) 1,000 [WLS17, Inf+21]
Movielens Adventure Fantasy Sci-Fi 2,008 (Men) 1,079 (Women) 1,000 [WLS17]
Book-Crossing Fiction Mystery Mystery 157 (Teenagers) 256 (Elderly) 100 [Inf+21, MK ̵14, CWM25]
Book-Crossing Fiction Fiction Adventure 380 (Teenagers) 102 (Elderly) 1,000 [CWM25, Dub+23]
Book-Crossing Fiction Adventure Adventure 316 (Teenagers) 40 (Elderly) 1,000 [Inf+21]
Alibaba H H H 1,096,218 (High Purchase Power) 306,722 (Low Purchase Power) $10^{6}$ [HND10, NFT19]
Alibaba L L L 166,714 (High Purchase Power) 2,432,915 (Low Purchase Power) $10^{7}$ [NFT19, Lu20]
Alibaba P H P 1,340,059 (High Purchase Power) 155,092 (Low Purchase Power) $10^{7}$ [MHV17, BW10]

### v2

@@ 5 String-Processing Applications
Using Theorem 2 on $\textsf{ST}(S)$, we can solve the Bottom-11 DR problem, which analogously to Bottom $k$-DR [sharma], asks for the string from $\mathcal{S}$ that contains the least number of occurrences of $P$, within the complexities of Theorem 3.
@@
Given a query with $P\in\Sigma^{m}$, we construct the suffix tree $\textsf{ST}(P)$ using Lemma 1, and compute $|\textsf{occ}_{P}(Q)|$, for every string $Q$ of length $q$ that occurs in $P$ in $\mathcal{O}(m)$ total time. We also use $\textsf{ST}(P)$ to mark every position $i^{\prime}$ of $P$ such that $P[i\mathinner{.\,.}i+q)=P[i^{\prime}\mathinner{.\,.}i^{\prime}+q)$, for some $i<i^{\prime}$, to avoid double-counting. We then spell $P$ in $\textsf{ST}(S)$ using suffix links [DBLP:books/cu/Gusfield1997], to iterate, for increasing $i$, over the loci (in $\textsf{ST}(S)$) of the longest prefixes of $P[i\mathinner{.\,.}i+q)$ that occur in $S$. We maintain a counter, which is initialized as zero, as follows. If $Q:=P[i\mathinner{.\,.}i+q)$ occurs in $\mathcal{S}$, $i$ is not marked, and $|\textsf{occ}_{P}(Q)|\in[f^{\min}_{v}-\varepsilon,f^{\max}_{v}+\varepsilon]$}, where $v$ is the node of $\textsf{ST}(S)$ with path-label $Q$ (if this is an implicit node, we take its nearest explicit descendant as $v$), we increase the counter by one. The value of the counter is output after processing all of $P$. Spelling $P$ takes $\mathcal{O}(m)$ time [DBLP:books/cu/Gusfield1997]; for every $q$-gram $Q$, we make $\mathcal{O}(1)$ elementary $\mathcal{O}(1)$-time operations, and hence the query time follows. ∎
@@ 6 Generalizations of SM
The $k$-RM problem asks for the $k$ most frequent colors in a range of an array [DBLP:journals/corr/abs-1101-4068]. Analogously, we define the $k$-SM problem asking for the $k$ most frequent colors in the leaves of the subtree rooted at a node of $\mathcal{T}$ (breaking ties arbitrarily). Corollary 2 is obtained by slightly modifying the SCM algorithm from Section 4.
@@
The SCM algorithm naturally extends for $k$-SM. For every node $v$ of $\mathcal{T}$, in Step 3, we now store the $k$ most frequent colors (breaking ties arbitrarily) and their frequencies as a list $(c^{1}_{v},f^{1}_{v}),\dots,(c^{k}_{v},f^{k}_{v})$. (If, for some node, we have fewer than $k$ colors, we leave some entries undefined.) We thus only need to verify that given $n>k$ integers we can select the $k$ largest ones in $\mathcal{O}(n)$ time to obtain the claimed generalization. We do this using the classic linear-time selection algorithm [DBLP:journals/jcss/BlumFPRT73]. After computing one list $(c_{1},f_{1}),\dots,(c_{k},f_{k})$ per node, we sort the lists by frequency using a single global radix sort. This takes $\mathcal{O}(kN)$ time since each frequency is at most $N$. ∎
@@ 7 Lower Bounds for Descendant-Mode in DAGs
Theorem 7 implies that an algorithm for DM that is similarly fast to our SCM algorithm is highly unlikely, as $p(N)=\mathcal{O}(N)$ and $q(N)=\mathcal{O}(1)$ for SM, and this would contradict the combinatorial BMM conjecture [AbboudFKLM24]. More generally, for any $\varepsilon>0$, there is no data structure for DM with construction time $\mathcal{O}(N^{\omega/2-\varepsilon})$ and sub-polynomial query time (unless the square matrix multiplication exponent $\omega$ is $2$). Under the combinatorial BMM conjecture, for any $\varepsilon_{1},\varepsilon_{2}\geq 0$, there is no combinatorial data structure for DM with construction time $\mathcal{O}(N^{3/2-\varepsilon_{1}})$ and query time $\mathcal{O}(N^{1/2-\varepsilon_{2}})$. These conditional lower bounds are identical to the ones for RM proved by Chan et al. [DBLP:journals/corr/abs-1101-4068].
@@ 8 Related Work
Krizanc et al. [DBLP:journals/njc/KrizancMS05] proposed a data structure for RM with $\mathcal{O}(1)$-time queries and $\mathcal{O}(N^{2}\log\log N/\log N)$ space, and another with $\mathcal{O}(\sqrt{N}\log N)$-time queries and $\mathcal{O}(N)$ space. Chan et al. [DBLP:journals/corr/abs-1101-4068] improved the time/space trade-off of the data structures of [DBLP:journals/njc/KrizancMS05] by designing an $\mathcal{O}(N+s^{2}/w)$-space data structure with $\mathcal{O}(N/s)$ query time, for any $s\in[1,N]$, where $w=\Omega(\log N)$ is the machine word size [DBLP:journals/corr/abs-1101-4068, Section 6]. By setting $s:=\lceil\sqrt{Nw}\rceil$, said data structure takes $\mathcal{O}(N)$ space and has $\mathcal{O}(\sqrt{N/w})$-time queries. In Baseline 2, we employ the data structure in [DBLP:journals/corr/abs-1101-4068, Section 3], as it is less expensive to construct [DBLP:journals/corr/abs-1101-4068, DBLP:journals/information/KarrasTKK24] and thus leads to a faster baseline for SM (Proposition 2). Greve et al. [DBLP:conf/icalp/GreveJLT10] showed a lower bound for RM: any data structure that uses $N\log^{\mathcal{O}(1)}N$ space needs $\Omega(\log N/\log\log N)$ time to answer an RM query, and any data structure that supports RM queries in $\mathcal{O}(1)$ time needs $N^{1+\Omega(1)}$ space. The current best conditional lower bound [DBLP:journals/corr/abs-1101-4068] indicates that answering $N$ RM queries on an array of size $\mathcal{O}(N)$ cannot be performed in $\mathcal{O}(N^{\omega/2-\varepsilon})$ time for any $\varepsilon>0$, where $\omega<2.3716$ [DBLP:conf/soda/WilliamsXXZ24] is the square matrix multiplication exponent. Dynamic and multidimensional versions of RM were studied in [DBLP:conf/icalp/SandlundX20, DBLP:conf/esa/El-Zein0MS18] and [DBLP:journals/tcs/DurocherEMT15], respectively. RM on paths of a tree (instead of subtrees as in SM) was studied in [DBLP:journals/algorithmica/DurocherSST16, DBLP:conf/esa/Gao022].
@@
SM can be used to construct a data structure for the $1$-DR problem; see Section 5. The $k$-DR problem has been studied in theoretical computer science; [DBLP:conf/soda/NavarroN12, DBLP:journals/jacm/HonSTV14, DBLP:journals/siamcomp/NavarroN17, DBLP:conf/soda/0001N25] proposed linear-space data structures with optimal or near-optimal query time but no efficient algorithms to construct them. Thus, these works are of theoretical interest, and the data structure we propose is the first practical approach for $k$-DR. $k$-DR differs from Top-$k$ document retrieval based on known patterns (keywords), for which we refer to [DBLP:conf/sigir/KhattabHE20, DBLP:conf/sigir/GouMLSS25].
@@
The UPM problem in Section 5 falls into the area of pattern mining [DBLP:books/sp/fpm14/Aggarwal14]. Specifically, it is somewhat related to discriminative (a.k.a emerging or contrast sets) pattern mining [DBLP:conf/kdd/DongL99, DBLP:journals/kais/DongL05, DBLP:journals/datamine/BayP01]. Given a collection of records (sequences [DBLP:journals/bib/LiuWGWH15, DBLP:journals/kais/MathonatNBK21, DBLP:conf/dasfaa/ChanKYT03], transactions [DBLP:journals/kais/DongL05], relational tuples [DBLP:journals/kais/DongL05], or vectors [DBLP:journals/datamine/BayP01]), the latter problem asks for mining all patterns (subsequences [DBLP:journals/bib/LiuWGWH15, DBLP:journals/kais/MathonatNBK21], substrings [DBLP:conf/dasfaa/ChanKYT03], itemsets [DBLP:journals/kais/DongL05], sets of relational attribute values [DBLP:journals/kais/DongL05], or sets of attribute/value pairs [DBLP:journals/datamine/BayP01]) which occur “disproportionately” in two or more parts of the collection that have different class labels. The disproportionality is measured based on difference in frequency [DBLP:journals/datamine/BayP01], growth rate [DBLP:conf/kdd/DongL99, DBLP:journals/kais/DongL05], weighted relative accuracy [DBLP:journals/kais/MathonatNBK21], or other measures [DBLP:journals/bib/LiuWGWH15]. The UPM problem is also relevant to fair pattern mining [DBLP:conf/sac/HajianMPDG14]. The latter problem asks for mining itemsets in a transaction dataset that do not produce association rules which are unprotected according to legally-grounded fairness measures [DBLP:conf/sac/HajianMPDG14]. None of the aforementioned approaches can solve UPM.
@@
Sequence comparison by means of $q$-grams is ubiquitous in bioinformatics [DBLP:journals/tcs/Ukkonen92]. It offers a faster alternative to using more expensive string measures such as edit distance. For instance, BLAST [Altschul1990] and FASTA [Pearson1988], two of the most widely-used tools for sequence-to-database search, are based on the notion of $q$-grams to report the best hits for a query.
@@
There are many other problems on leaf-colored [DBLP:conf/cpm/Hui92, steel, DBLP:conf/soda/Muthukrishnan02] and node-colored [pawel, shlomo, golovach] trees.
@@ 9 Experimental Evaluation
We used $4$ benchmark datasets (see Table 2): (1) WebKB [craven1998learning], which is a collection of webpages of computer science departments of various universities; (2) Genes, which is a collection of DNA sequences between two markers flanking the human X chromosome centromere [sayers2023database]; (3) News [zhou2015pattern], which is a collection of documents from the Newsgroups dataset, and (4) Vir [ncbi_genome_2025], which is a collection of viral genomes. As the baselines could not run on large datasets, we applied them to samples of the first two datasets, constructed by selecting strings that have the same length uniformly at random; see Table 2.
@@
WebKB [craven1998learning] Web 26 790,340 32.98 26,068,175 38,491,476
Genes [ncbi_handbook_chapter6] Biology 4 800,000 1,250.01 1,000,007,888 2,030,921,400
News [zhou2015pattern] News 26 4,781 725.556 3,768,883 5,293,347
Vir [ncbi_genome_2025] Biology 4 143,588 29,085 4,176,208,246 7,313,326,212
@@
WebKB-sam [craven1998learning] Web 26 31,030 100 3,103,000 4,876,306
Genes-sam [ncbi_handbook_chapter6] Biology 4 10,000 200 2,000,000 3,753,134
@@
We examined the impact of the two problem parameters, $N$ and $\Delta$, on runtime and space. We also showcase the benefit of our approach in UPM, CQS, $k$-DR, and phylogenetic tree annotation. In these applications, we used four real datasets.
@@
Figs. 7(a) to 7(c) show the runtime for varying $N$ and fixed $\Delta$. Our SCM algorithm was faster than the fastest baseline, BA3, by at least one order of magnitude and $24$ times on average. BA3 in turn was faster than both BA1 and BA2 by at least two orders of magnitude on average. All algorithms scaled linearly with $N$, in line with their time complexities, except BA2 whose time complexity is $\mathcal{O}(N\sqrt{N})$. In fact, since the space complexity of BA2 is also $\mathcal{O}(N\sqrt{N})$, it did not terminate for strings with more than 1 million letters; it needed more than the 1TB of memory that was available. Figs. 7(d) to 7(f) show the peak memory consumption for the experiments of Figs. 7(a) to 7(c). SCM needed on average $26\%$ and up to $58\%$ less memory compared to the best baseline, BA3, as storing counts for the single-color trees needs less memory than merging trees. BA3 in turn needed two orders of magnitude less memory on average compared to both BA1 and BA2, as its space complexity is $\mathcal{O}(N)$, while that of BA1 and BA2 is $\mathcal{O}(N\Delta)$ and $\mathcal{O}(N\sqrt{N})$, respectively.
@@
Figs. 8(a) to 8(c) show the runtime for varying $\Delta$ and fixed $N$; we fixed $N=300,000$ for the WebKB-sam dataset, $N=162,000$ for News, and $N=200,000$ for Genes-sam by taking the prefix of length $\lfloor N/\Delta\rfloor$ of each string. Our SCM algorithm was faster than the fastest baseline, BA3, by at least an order of magnitude and $21$ times on average and, as expected by its time complexity, its runtime was not affected by $\Delta$. BA3 in turn was faster than both BA1 and BA2 by two orders of magnitude on average, which is in line with the time complexities of these algorithms. Figs. 8(d) to 8(f) show the peak memory consumption for the experiments of Figs. 8(a) to 8(c). SCM needed at least $20\%$ and up to $50\%$ less memory compared to the best baseline, BA3, for the same reason as in the experiments of Figs. 7(d) to 7(f). BA3 in turn was more space-efficient than both BA1 and BA2 by more than two orders of magnitude on average, which is in line with the space complexities of these algorithms. Also, BA2 uses more memory as $\Delta$ increases, as the string index it uses gets larger.
@@
Figs. 9(a) to 9(c) show the runtime for varying $N$ and fixed $\Delta$. Both SCM and BA3 scaled linearly with $N$, in line with their time complexities, but SCM was at least $18$ and up to $40$ times faster, as BA3 has an extra $\log\Delta$ term in its time complexity which is $17$ to $20$ depending on the dataset. SCM is practical; it took less than $20$ minutes when applied to the entire Vir dataset whose total length is $4.2$ billion letters (and suffix tree has over $7.3$ billion nodes). Figs. 9(d) to 9(f) show the peak memory consumption for the experiments of Figs. 9(a) to 9(c) . SCM needed $15\%$ less memory compared to BA3 on average, which is in line with the space complexities of the algorithms. This shows again the benefit of the count merging in SCM compared to the tree merging in BA3.
@@
Figs. 9(g) to 9(i) show the runtime for varying $\Delta$ and fixed $N$; we fixed $N$ by removing letters from each string evenly, so that the remaining strings have total length $N=2,000,000$ for WebKB, $N=62,500,000$ for Genes, and $N=10^{9}$ for Vir. Again, SCM was faster than BA3 by at least $15$ and $26$ times on average. For example, when $\Delta=800,000$ in Genes (see Fig. 9(h)), SCM took only about $20$ seconds while BA3 about $8$ minutes. Figs. 9(j) to 9(l) show the peak memory consumption for the experiments of Figs. 9(g) to 9(i). Again, SCM was more space-efficient than BA3; it needed at least $9\%$ and up to $17\%$ less memory.
@@
We consider collections whose strings represent different user subpopulations and solve the UPM problem with large $\varepsilon$. As will be discussed later, each letter in each string represents a movie genre, book genre, or product price tier, and UPM discovers all substrings with bounded difference in terms of their frequencies in the strings of the collection. There is no other restriction to the type of patterns that can be discovered, which makes UPM challenging. For example, if one could restrict the length of the mined patterns and/or their items (e.g., to be semantically similar movie genres), then the problem space becomes smaller and other approaches such as those based on inverted indexes may be more suitable. However, actionable patterns (e.g., the pattern comprised of the $4$ non-semantically similar letters/items Comedy Drama Romance Comedy that we mined) may be missed. In our case study, we discover patterns revealing behavioral preferences that prevail in these subpopulations. For each pattern, we refer to literature supporting that it is in fact prevailing in these subpopulations.
@@
We processed three datasets: Movielens [harper2015movielens], Book-Crossing [ziegler2005improving], and Alibaba [pei2019value]. The processed datasets are comprised of two strings each, and they can be found at our code and data link. Each string corresponds to a different user subpopulation, and in each string each distinct letter corresponds to a distinct entity (movie genre, book genre, price tier). In Movielens, one string corresponds to $1,709$ men and the other to $1,709$ women. Each string has length $4,000,033$ and contains genres of rated movies, ordered chronologically. The total number of distinct genres (alphabet size) is $18$. In Book-Crossing, one string corresponds to $9,164$ teenagers (1010-$18$ years old) and the other to $4,252$ elderly (6565-$100$ years old). Each string has length $14,145$ and contains book genres, ordered by users’ ratings for book genres from high to low. The genres were found by mapping book ISBNs to genres using ChatGPT-3.5-turbo [openai2023chatgpt], which is remarkably good in this task [raj]. The total number of distinct book genres (alphabet size) is $10$. In Alibaba, one string corresponds to $928,622$ users with high purchase power ($\geq 15$) and the other to $894,770$ users with low purchase power (<15<15). Each string has length $34,651,424$ and contains price tiers of products, ordered in the way users browsed them. The total number of distinct price tiers (alphabet size) is $4$. Specifically, the letters L, M, H, and P, correspond to the Low, Middle, High, and Premium price tier, respectively.
@@
Movielens Action Drama War Action 1,551 (Men) 588 (Women) 1,000 [wuhr2017tears, krcmar2005uses]
Movielens Comedy Drama Romance Comedy 899 (Men) 1,860 (Women) 1,000 [wuhr2017tears, infortuna2021inner]
Movielens Adventure Fantasy Sci-Fi 2,008 (Men) 1,079 (Women) 1,000 [wuhr2017tears]
Book-Crossing Fiction Mystery Mystery 157 (Teenagers) 256 (Elderly) 100 [infortuna2021inner, awriterofhistory_2014, currie2025reading]
Book-Crossing Fiction Fiction Adventure 380 (Teenagers) 102 (Elderly) 1,000 [currie2025reading, dubourg2023exploratory]
Book-Crossing Fiction Adventure Adventure 316 (Teenagers) 40 (Elderly) 1,000 [infortuna2021inner]
Alibaba H H H 1,096,218 (High Purchase Power) 306,722 (Low Purchase Power) $10^{6}$ [han2010signaling, ngwe2019impact]
Alibaba L L L 166,714 (High Purchase Power) 2,432,915 (Low Purchase Power) $10^{7}$ [ngwe2019impact, lu2020status]
Alibaba P H P 1,340,059 (High Purchase Power) 155,092 (Low Purchase Power) $10^{7}$ [mundel2017exploratory, berger2010subtle]